A 1500 kg VW is heading 40 m/s in a straight line. A 4000 kg Cadillac is heading
directly for it at 60 m/s. Find the velocity (magnitude and direction) of the center of
mass.
1
Expert's answer
2020-12-10T11:09:55-0500
"V=\frac{m_2v_2-m_1v_1}{m_1+m_2}\\V=\frac{(4000)(60)-(1500)(40)}{1500+4000}=32.7\frac{m}{s}"


Direction is direction of motion of cadillac.